# ExPath: Targeted Pathway Inference for Biological Knowledge Bases via Graph Learning and Explanation

### v1

@@
Abstract.
Biological knowledge bases provide systemically functional pathways of cells or organisms in terms of molecular interaction. However, recognizing more targeted pathways, particularly when incorporating wet-lab experimental data, remains challenging and typically requires downstream biological analyses and expertise. In this paper, we frame this challenge as a solvable graph learning and explaining task and propose a novel pathway inference framework, ExPath, that explicitly integrates experimental data, specifically amino acid sequences (AA-seqs), to classify various graphs (bio-networks) in biological databases. The links (representing pathways) that contribute more to classification can be considered as targeted pathways. Technically, ExPath comprises three components: (1) a large protein language model (pLM) that encodes and embeds AA-seqs into graph, overcoming traditional obstacles in processing AA-seq data, such as BLAST; (2) PathMamba, a hybrid architecture combining graph neural networks (GNNs) with state-space sequence modeling (Mamba) to capture both local interactions and global pathway-level dependencies; and (3) PathExplainer, a subgraph learning module that identifies functionally critical nodes and edges through trainable pathway masks. We also propose ML-oriented biological evaluations and a new metric. The experiments involving 301 bio-networks evaluations demonstrate that pathways inferred by ExPath maintain biological meaningfulness. We will publicly release curated 301 bio-network data soon. The source code is available at: https://anonymous.4open.science/r/ExPath
pathway, graph, SSM, explainable, pLM, biology
copyright: acmcopyrightjournalyear: 2025doi: 10.1145/1122445.1122456conference: The 31th ACM SIGKDD Conference on Knowledge Discovery and Data Mining; August 3–7, 2025; Toronto, Canadabooktitle: Proceedings of the 31th ACM SIGKDD Conference on Knowledge Discovery and Data Mining, August 3–7, 2025, Washington, DC, USAisbn: 978-1-4503-XXXX-X/18/06journalyear: 2025copyright: acmcopyrightconference: 2025 ACM SIGKDD Conference on Knowledge Discovery and Data Mining; Toronto, Canada; ccs: Applied computing Bioinformaticsccs: Computing methodologies Machine learning approaches
1. Introduction
Decades of research have generated extensive network biology data, revealing that systems, from cells to organisms, can be considered molecular networks and have been reported in extensive literature (Barabasi and Oltvai, 2004; Alm and Arkin, 2003; Ideker and Krogan, 2012). These data resources have been compiled into functional biological knowledge bases, such as KEGG (Kanehisa and Goto, 2000) and STRING (Szklarczyk et al., 2023), framing bio-network resources (Karlebach and Shamir, 2008; Phizicky and Fields, 1995) that document various interactions between molecules (e.g., genes or proteins) to describe how molecular behaviors relate to biological systems. These knowledge bases are now widely used to mine and interpret wet-lab experimental data, enabling researchers to study disease mechanisms (Barabási et al., 2011), drug interactions (Zhang et al., 2017), and potential therapeutic targets (Li et al., 2021).
@@ 1. Introduction
Figure 1. Biological knowledge bases lack specificity when integrating experimental data. This example illustrates two experimental datasets with different mutations that share the same disease pathway structure, yet fail to reveal the distinct interactions responsible for these differences.
While biological knowledge bases are comprehensive and continuously updated, a main concern remains: they lack specificity for experimental data. The bases cannot provide the information on which interactions are more relevant to the given data, even though their main purpose is to interpret the data. As shown in Figure 1, two experimental datasets from the same disease with different mutations share the same network structure, but fail to reveal which distinct interactions account for these differences. However, inferring key molecular interactions is crucial for understanding the potential roles of genes or proteins, potentially accelerating new biomarker discovery (Zitnik et al., 2024). Conceptually, in-lab experimental data, such as amino acid sequences or gene expression profiles, is typically generated under specific research objectives or experiments, often focusing on proteins or genes of interest. However, knowledge bases provide only generic networks. They collect “meaningful” molecular interactions via automated text mining and manual curation (Kanehisa et al., 2024), but do not incorporate the specific data that “confer such meaningfulness” into the network-building process. As a result, they cannot provide data-specific interactions, and this generalization can lead to misinterpretations of experimental results. A few complementary bio-tools, such as BLAST (Altschul et al., 1997), combined with downstream analytic methods (Huang et al., 2009), can help infer bio-networks. However, these tools are not user-friendly for non-experts, as they require domain expertise for implementation and information management (Madden, 2013). Moreover, the downstream analysis generally provides empirical explanations for the data, and sometimes, extensive in-lab evaluation, such as pairwise examinations among hundreds to thousands of genes (Zaman et al., 2013), is needed. There is thus an urgent need for a method that integrates bio-network resources with experimental data to infer targeted interactions, thereby facilitating the efficient use of knowledge bases.
Bio-network inference, including computational and learning-based methods, has emerged as a promising solution for mining targeted interactions. Such bio-network inference has attracted great attention from both computational biology and machine learning researchers (Wu et al., 2021; Kang et al., 2022). Existing methods typically form bio-network topology as graph data, embed experimental data as node features, and define proper objective functions to infer meaningful subgraphs. Computational methods typically incorporate statistical node-centric metrics in graph theory, such as node degrees (Karger, 1994; Kashtan et al., 2004a), centrality (Haynes et al., 2013; Wang et al., 2014), betweenness (Nacher and Akutsu, 2016), or PageRank scores (Iván and Grolmusz, 2011; Maehara et al., 2014), to evaluate the importance of nodes within in a graph. The edges connecting these top-ranked nodes, or those highlighted during node evaluations, can be identified as more targeted interactions. However, such objectives lack explicit inference of interactions, i.e., edges, and are often computationally intractable for large bio-networks (Pavlopoulos et al., 2011). Moreover, computational methods cannot fully exploit experimental data, as the statistical metrics do not integrate node features into scoring schemes. By contrast, machine learning methods, particularly graph neural networks (GNNs), model both network topology and node-level attributes for subgraph inference in a data-driven manner (Wang et al., 2021; Ravindra et al., 2020). These methods set up objective functions such as link prediction and graph reconstruction to explicitly infer targeted interactions. More importantly, experimental data can directly influence the objective functions through node aggregation in GNN learning, ensuring that the model outputs are more specific to the data. However, most existing works are task-specific, and their objectives aim to learn the general graph structure accurately, including irrelevant interactions (Wu et al., 2021; Zhao et al., 2021). Some works propose to gradually infer subgraph structure, mitigating the constraints of prior general bio-network information (Kang et al., 2022; Li et al., 2024). Nonetheless, they do not explicitly learn the distinct interactions unique to different experimental data. Furthermore, existing methods typically require downstream biological analysis to interpret the model outputs or interactions. We acknowledge that current explorations in bio-network inference remain nascent.
In this paper, we study a novel and critical problem of developing a bio-network inference framework that explicitly generates data-specific interactions while maintaining biological plausibility. We note that this is a non-trivial task. In essence, we mainly face three challenges. $\rm(\hskip 1.79993pt1\hskip 1.79993pt)$ Qualitative interaction inference objective. The first challenge involves formulating a new objective that not only learns but also qualitatively assesses interactions directly. The model outputs, therefore, enable the direct interpretation of targeted interactions, eliminating the need for downstream analysis. $\rm(\hskip 1.79993pt2\hskip 1.79993pt)$ Pathway modeling. One primary focus of knowledge bases is on “pathways” (Kanehisa et al., 2024), i.e., connected multi-interactions, representing a sequence of events where one protein interaction triggers the next, ultimately leading to a defined outcome. However, existing works often treat all interactions equally and uniformly, focusing primarily on isolated interactions. A reliable method should explicitly incorporate pathway information into the modeling process to maintain biological plausibility. $\rm(\hskip 1.79993pt3\hskip 1.79993pt)$ ML-oriented evaluation. Currently, there is no standardized quantitative evaluation framework tailored for machine learning models. Most evaluations depend on qualitative methods, such as enrichment analysis (Huang et al., 2009), to determine whether the resulting interactions are biologically relevant, which requires domain expertise. A quantitative evaluation method, designed to directly assess the model outputs, would bridge the gap between computational and biological disciplines.
In this paper, we present ExPath, a deep learning framework for inferring targeted pathways for bio-networks. To tackle the above challenges, $\rm(\hskip 1.79993pt1\hskip 1.79993pt)$ we formulate bio-network inference as a subgraph learning and explanation task. Subgraphs, contributing most significantly to the learning objective, can be identified as targeted interactions. $\rm(\hskip 1.79993pt2\hskip 1.79993pt)$ To ensure these subgraphs capture high-order pathways, technically, we propose two novel models: PathMamba, a hybrid learning model, combines graph neural networks (GNNs) with state-space sequence modeling (Mamba) to learn both local interactions and global pathway-level dependencies; PathExplainer, a novel variation of GNNExplainer, designed as a subgraph explanation module, identifies objective-critical pathways. We take amino acid (AA) sequences as the reference experimental data since many biological databases organize pathway information at the protein level (Finn et al., 2014). We utilize a large protein language model (pLM) (ESM-2 (Lin et al., 2023)) to encode AA sequences into graph embeddings, driving ExPath to mine targeted pathways effectively. $\rm(\hskip 1.79993pt3\hskip 1.79993pt)$ We propose an evaluation workflow that directly incorporates model-derived weights of subgraphs to quantitatively assess their biological significance. Overall, our contributions are:
  • Formulating Bio-network Inference Problem. We formulate and make an initial investigation on a novel research problem of inferring data-specific pathways for bio-networks.
  • Proposing New Framework. ExPath consists of PathMamba and PathExplainer, tailored for pathway-level modeling, can interpret data directly. It achieves the best fidelity+ and fidelity- across 10 baselines, showing both sufficiency and necessity.
  • Impacting Biological Relevance. We propose ML-oriented biological evaluations and a new metric. The experiments involving 301 bio-networks evaluations demonstrate that pathways inferred by ExPath maintain biological meaningfulness.
  • Datasets. We collected all available human networks from the KEGG, constructed machine-learning-ready datasets, and will release them soon.
2. Preliminary and Problem Setting
@@
Figure 2. Overview of ExPath. ExPath comprises three components. (1) Data encoding with a large protein language model and RWPE for node attributes (AA sequences) learning; (2) PathMamba combining graph neural networks with state-space sequence modeling (Mamba) to capture both local interactions and global pathway-level dependencies for pathway information learning; and (3) PathExplainer identifies functionally critical nodes and edges through trainable pathway masks for targeted pathway inference.
2.1. Pathways in Biological Knowledge Bases
Knowledge bases organize these pathways as interconnected graphs, where nodes represent proteins and edges denote interactions (e.g., enzymatic reactions, regulatory effects). These pathways help researchers understand how different molecules work together rather than looking at just one interaction at a time. For example, while protein-protein interactions (PPI) (Phizicky and Fields, 1995) capture pair-wise molecular associations, pathways show how many proteins are linked in complex networks. This helps scientists see the “big picture,” revealing how different parts of a cell cooperate to perform important tasks (Kanehisa et al., 2024). Learning from amino acid (AA) sequence data is challenging due to its inherent complexity. Even slight variations can lead to significant structural changes, potentially disrupting protein functionality within pathways. Several studies focus on extracting meaningful features from AA sequences, like like AlphaFold (Jumper et al., 2021). In this paper, we investigate the mapping of AA sequence data to corresponding pathway bio-networks.
2.2. Problem Setting
Definition 1 (Amino Acid sequence data). Consider a collection of $M$ AA sequences, denoted as $\mathbf{S}=\left\{S^{(m)}\right\}_{m=1}^{M}$, where each sequence $S^{(m)}$ has a length $L^{(m)}$ (which may vary across $m$) and is represented as $S^{(m)}=\left[s_{1}^{(m)},s_{2}^{(m)},\dots,s_{L^{(m)}}^{(m)}\right]$. Here, each amino acid $s_{i}^{(m)}$ belongs to the standard set of $20$ canonical amino acids.
@@ 2.2. Problem Setting
Definition 2 (Knowledge bio-networks). The bio-networks can be represented as a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ denotes the vertices (e.g., AA sequences) and $\mathcal{E}$ is the set of edges, representing molecular interactions. Let $\mathbf{G}={\{\mathcal{G}^{(m)}\}}_{m=1}^{M}$ denote a dataset comprising $M$ bio-networks. Each $\mathcal{G}^{(m)}$ is associated with a label $y^{(m)}\in\mathbf{Y}$, indicating its primary biologically functional class (e.g., metabolism, genetic information processing, and human diseases).
Problem (Targeted Pathway Inference). We frame this research problem as a two-phase graph learning and subgraph explanation task. Given $\mathbf{G}$ and the labels $\mathbf{Y}$, we focus on two main tasks:
  • Task-1 (Classification): Train a classifier $F(\cdot)$ to predict the function label $y^{(m)}$ of an unseen bio-network $\mathcal{G}^{(m)}$, driven by the node features, i.e., AA sequence data $S$. To perform accurate classification, it can effectively learn various pathways in biological knowledge bases.
  • Task-2 (Explanation): For each predicted function class $y\in\mathbf{Y}$, employ an explainer $E(\cdot)$ to extract a class-specific subgraph $\hat{\mathcal{G}}\subseteq\mathcal{G}$ that highlights the most influential pathways contributing to the classification outcome.
In plain words, a subgraph learned from various biological pathways and capable of accurately predicting its associated bio-network can be considered as representing targeted pathways. Let $\hat{\mathcal{G}}_{y}=E(\mathcal{G},y;S)$ denote subgraphs for bio-network $y$, integrated to experimental data $S$. Here, we hold three hypotheses for our method:
  • Hypothesis-1. The classifier $F(\cdot)$ leverages bio-network topology and AA seq features to achieve high prediction accuracy while maintaining balanced performance across all classes.
  • Hypothesis-2. The explainer $E(\cdot)$ infers subgraphs that retain high fidelity for class discrimination and exhibit distinct structures across classes, reflecting unique biological mechanisms.
  • Hypothesis-3. Both $F(\cdot)$ and $E(\cdot)$ must maintain biologically plausible mechanisms, ensuring that the inferred subgraphs capture high-order pathways and hold biological meaningfulness.
3. Proposed method
3.1. Framework
@@
ExPath comprises three components: large protein model encodings, graph-based classification, and post-hoc subgraph explanation, as shown in Figure 2. First, we encode AA sequence data using a pre-trained large protein model, ESM-2 (Lin et al., 2023) (see Section 3.2 ), to serve as node attributes. To address (Task-1, Hypothesis-1), PathMamba, a classifier combining graph neural networks (GNNs) with state-space sequence modeling (Mamba), is to capture both local node-pair interactions and global pathway-level dependencies (see Section 3.3). To address (Task-2, Hypothesis-2), PathExplainer, an explainer method trained with pathway-wise masks (see Section 3.4), aims to identify the most influential subgraphs. We explicitly integrate pathway-level information into both models to satisfy Hypothesis-3. The large protein model ensures biologically meaningful protein representations, PathMamba leverages various pathway information in knowledge bases for robust classification, and PathExplainer highlights minimal subgraphs that drive the final predictions, offering interpretable insights into key pathways. We evaluate ExPath from both machine learning and biological perspectives, as detailed in Section 4.
3.2. Data Encoding for Node Attributes
3.2.1. Large Protein Language Model Encoding
The ESM-2 model (Lin et al., 2023), pre-trained on over 60 million AA sequences with parameter scaling up to 15 billion, is employed to encode our data. We evaluate different parameter variants of ESM-2, and the results are presented in Table 1. Formally, each $S^{(m)}$ is tokenized and passed through stacked Transformers. The output is a token vector, denoted as: $\mathbf{h}_{i}=\mathbf{H}_{1}^{(L)},$ where $\mathbf{H}_{1}^{(L)}\in\mathbb{R}^{d}$ is the embedding of the first token from the $L$-th (last) Transformer layer, serving as data representation.
3.2.2. Positional Encoding
To address a fundamental limitation of GNNs (Xu et al., 2019) or hybrid models (Rampášek et al., 2022) to distinguish nodes with identical local structures, we apply a random-walk-based positional encoding (RWPE) base on a diffusion process (Dwivedi et al., 2022), defined as: $p_{i}=[RW_{ii},RW_{ii}^{2},\cdots,RW_{ii}^{k}]\in\mathbb{R}^{k}$ where $RW=AD^{-1}$ is the random walk operator, constructed by the adjacency matrix $A$ and degree matrix $D$. For each node $i$, $RW_{ii}^{k}$ captures the probability of returning to node $i$ after $k$ steps of random walk.
The final node representation combines the sequence-level features from ESM-2 and the structural information from the graph. Specifically, the sequence embedding $\mathbf{h}_{i}$ and the positional encoding $\mathbf{p}_{i}$ are concatenated and passed through a linear layer to obtain the final representation: $\mathbf{x}_{i}=\text{Linear}([\mathbf{h}_{i}\|\mathbf{p}_{i}]),$ where $[\mathbf{h}_{i}\|\mathbf{p}_{i}]\in\mathbb{R}^{d+K}$ denotes the concatenation of $\mathbf{h}_{i}\in\mathbb{R}^{d}$ and $\mathbf{p}_{i}\in\mathbb{R}^{K}$. The linear transformation ensures dimensionality reduction and effective integration of both global protein sequence features and local graph structural information. This final node feature $\mathbf{x}_{i}\in\mathbb{R}^{d}$ is optimized for downstream tasks such as graph classification.
3.3. PathMamba: Pathway Information Learning
@@
PathMamba integrates the Graph Isomorphism Network (GIN) with a novel pathway-wise Mamba model. It leverages the strengths of both global selective modeling mechanisms and message-passing GNNs. Specifically, inspired by GPS (Rampášek et al., 2022), our model avoids early-stage information loss that could arise from using GNNs exclusively in the initial layers. We hence employ pathway-wise global aggregation in combination with an efficient Mamba mechanism (Gu and Dao, 2024). At each layer, node and edge features are updated by aggregating the outputs of a pathway-wise Mamba aggregation as:
(1) $X^{l+1},$ $=$ $\texttt{PathMamba}^{l}\left(X^{l},A\right),$
(2) computed $\textrm{computed as}\ \ \ X^{l+1}_{L},$ $=$ $\texttt{LocalGIN}^{l}\left(X^{l},A\right),$
(3) $X^{l+1}_{G},$ $=$ $\texttt{GlobalMamba}^{l}\left(X^{l},A\right),$
(4) $X^{l+1},$ $=$ $\texttt{MLP}^{l}\left(X^{l+1}_{L}+X^{l+1}_{G}\right),$
@@
3.3.1. Node-wise local aggregation
Node features are updated by aggregating information from their local neighbors. The GIN operation can be expressed as:
(5) $X^{l+1}_{L}=\text{ReLU}\left(W^{l}\cdot\big{(}(1+\epsilon)X^{l}+\sum_{j\in\mathcal{N}(i)}X^{l}_{j}\big{)}\right),$
where $\mathcal{N}(i)$ represents the set of neighbors of node $i$, $W^{l}$ is the learnable weight matrix at layer $l$, and $\epsilon$ is a trainable parameter controlling the importance of self-loops. This ensures a high level of expressivity for local feature aggregation.
3.3.2. Pathway-wise global aggregation
To capture long-range dependencies within pathways, we propose (1) random pathway sampling and (2) sequential pathway modeling in PathMamba.

### v2

# Targeted Pathway Inference for Biological Knowledge Bases via Graph Learning and Explanation

@@
Abstract
Retrieving targeted pathways in biological knowledge bases, particularly when incorporating wet-lab experimental data, remains a challenging task and often requires downstream analyses and specialized expertise. In this paper, we frame this challenge as a solvable graph learning and explaining task and propose a novel subgraph inference framework, ExPath, that explicitly integrates experimental data to classify various graphs (bio-networks) in biological databases. The links (representing pathways) that contribute more to classification can be considered as targeted pathways. Our framework can seamlessly integrate biological foundation models to encode the experimental molecular data. We propose ML-oriented biological evaluations and a new metric. The experiments involving 301 bio-networks evaluations demonstrate that pathways inferred by ExPath are biologically meaningful, achieving up to 4.5× higher Fidelity+ (necessity) and 14× lower Fidelity- (sufficiency) than explainer baselines, while preserving signaling chains up to 4× longer.
Introduction
Decades of research have revealed that systems, from cells to organisms, can be considered biological networks (Ideker and Krogan 2012). These networks have been compiled into public knowledge bases such as KEGG (Kanehisa and Sato 2020) and STRING (Szklarczyk et al. 2023), which document molecular (e.g., among genes or proteins) interactions and their roles in cellular functions. While knowledge bases are continuously updated, a primary concern remains: they lack specificity for experimental data. The main objective of biological knowledge bases is to cover all possible interactions in a system. These networks are general and static. In contrast, experimental studies focus on one specific condition or dataset, where only a subset of the network is actually relevant. Our objective is to identify which interactions are active, meaningful, or target-specific in the given data, as shown in Figure 1. In this paper, we propose to infer the bio-networks that capture targeted interactions from experimental data, thereby facilitating downstream analyses.
Many researchers have formulated this bio-network inference as a graph learning problem. In this setting, interactions in a bio-network are modeled as graph edges, and experimental data are embedded as node features. Various computational and machine learning methods have been proposed to infer meaningful targeted graph structures. Computational methods often rely on statistical node-centric metrics (Nacher and Akutsu 2016) to evaluate the importance of nodes. Edges connected to highly ranked nodes are considered more important. However, such objectives lack explicit inference of interactions and are computationally intractable for large bio-networks (Yang et al. 2024). Machine learning methods, particularly graph neural networks (GNNs), define explicit objectives such as link prediction or graph reconstruction, enabling direct inference of network structure (Ravindra et al. 2020). Importantly, experimental data influence the learning process via node feature aggregation, making the inferred interactions more specific to the dataset. However, existing methods are still in the early stages of exploration, and several key limitations remain unaddressed.
@@ Introduction
Figure 1: This example illustrates two experimental datasets with different mutations (red and blue) that are mapped onto the same disease network, yet fail to reveal the distinct interactions that account for their differences.
  • Implicit targeted interaction inference. Their objectives aim to reconstruct the general graph structure accurately, including irrelevant interactions. Some works propose to gradually infer subgraph structure, weakening the influence of prior general bio-network information (Li et al. 2024). However, they still fail to explicitly identify the distinct interactions unique to different experimental data.
  • Lacking pathway modeling. Existing works treat all interactions equally and independently, overlooking long-range dependencies in biological pathways. In reality, biological systems typically exhibit multi-step interactions, where one protein interaction triggers another, eventually leading to specific cellular outcomes.
  • Inadequate biologically plausible evaluation. Existing methods typically require downstream biological analysis to qualitatively interpret the inferred interactions, which requires domain expertise. There is a lack of quantitative evaluation methods tailored for machine learning models.
To tackle the above challenges, we propose ExPath, a deep learning framework for inferring targeted data-specific pathways in bio-networks, with the following contributions:
  • Graph explanation formulation for explicit interaction inference. we formulate bio-network inference as a subgraph learning and explanation task, and hence propose a graph-based model equipped with GNNExplainer. Subgraphs, contributing most significantly to the learning objective, are explicitly identified as targeted interactions.
  • Pathway-level encoding and explaining. To ensure these subgraphs capture high-order pathways, technically, we propose two novel models: PathMamba, a hybrid learning model, combines GNNs with state-space sequence modeling (Mamba) to learn both local interactions and global pathway-level dependencies; PathExplainer identifies objective-critical pathways by learning novel pathway masks. We also provide a theoretical analysis of ExPath’s expressiveness and show that identified pathways capture higher-order structural patterns.
  • A novel ML-oriented biological evaluation. We propose an evaluation workflow that directly incorporates model-derived subgraph importance scores to quantitatively assess their biological relevance.
ExPath can seamlessly integrate biological foundation models, and in this work, we use the large protein language model, ESM-2 (Lin et al. 2023), as a case encoder. We collect all available human pathway networks from KEGG (Kanehisa et al. 2024), resulting in 301 bio-network, using amino acid (AA) sequences as reference experimental data. Extensive experiments demonstrate that the pathways inferred by ExPath are biologically meaningful, achieving up to 4.5× higher Fidelity+ (necessity) and 14× lower Fidelity- (sufficiency) than explainer baselines, while preserving signaling chains up to 4× longer.
Related Work
Existing methods can be grouped into statistical topology-driven and data-driven deep graph learning methods.
- Topology-driven Methods. They use statistical metrics on structural properties of graphs, such as node degrees (Karger 1994; Kashtan et al. 2004a), centrality (Haynes, Hedetniemi, and Slater 2013; Wang et al. 2014; Nacher and Akutsu 2016), betweenness, or PageRank scores (Iván and Grolmusz 2011; Maehara et al. 2014) to infer which substructures exhibit a more significant influence on the overall topology, thereby identifying more targeted interactions.
- Deep Graph Learning Methods. They incorporate experimental data during the learning process by embedding data as node representations. They train GNNs with suitable objectives, such as link prediction or graph reconstruction (Yue et al. 2020; Zhang et al. 2021; Muzio, O’Bray, and Borgwardt 2021), and the links that contribute most to these objectives can be considered the targeted interactions. For instance, the works of (Hamilton, Ying, and Leskovec 2017; Gligorijević et al. 2021; Cheng et al. 2021) have been validated to predict protein functions within protein-protein interaction (PPI) networks. Moreover, GNN models have been applied to incorporate RNA-Seq data for tasks like predicting disease states and cell-cell relationships (Wang et al. 2021; Ravindra et al. 2020).
Limitations of Previous Work. The topology-driven methods focus only on graph edges. They cannot incorporate experimental data to infer biological networks. While GNN-based methods can generate targeted interactions in a data-driven manner, their objectives do not explicitly focus on inferring networks and are typically task-specific. In contrast, our method focuses on directly explaining graph representations of bio-networks under specific experimental data.
Problem Formulation
@@
Figure 2: Overview of ExPath. Our method comprises two novel components. (1) PathMamba combining graph neural networks with state-space sequence modeling (Mamba) to capture both local interactions and global pathway-level dependencies for pathway information learning; and (2) PathExplainer identifies functionally critical nodes and edges through trainable pathway masks for targeted pathway inference.
@@
Definition 1 (Knowledge bio-networks). The bio-networks can be represented as a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ denotes the vertices, each representing a molecule such as a gene or protein, and $\mathcal{E}$ is the set of edges, representing molecular interactions. Let $\mathbf{G}={\{\mathcal{G}^{(m)}\}}_{m=1}^{M}$ denote a dataset comprising $M$ bio-networks. Each $\mathcal{G}^{(m)}$ is associated with a label $y^{(m)}\in\mathbf{Y}$, indicating its primary biologically functional class such as metabolism or human diseases.
Definition 2 (Molecular experimental data). For each node $v\!\in\!\mathcal{V}$, we are given a condition-specific feature vector $\mathbf{x}_{v}\!\in\!\mathbb{R}^{d}$ derived from molecular experiments (e.g., amino-acid sequence embeddings, gene-expression counts, or protein abundances ). Collecting all nodes yields the matrix $\mathbf{X}^{(m)}=\!\bigl{[}\mathbf{x}_{v}\bigr{]}_{v\in\mathcal{V}^{(m)}}$ for network $\mathcal{G}^{(m)}$.
Problem 1 (Bio-network inference). Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ be a curated graph and the node features $\mathbf{X}=\{\mathbf{x}_{v}\}_{v\in\mathcal{V}}$ obtained from condition-specific experimental data (e.g., amino-acid embeddings). Although $\mathcal{G}$ is static, the pair $(\mathcal{G},\mathbf{X})$ constitutes a data-specific graph that reflects the molecular state of the same pathway under the given experiment. To this end, we formulate this problem as a two-stage sub-graph learning and explaining task.
  • Task-1: Graph representation learning and classification.  Learn a classifier $F(\mathcal{G},\mathbf{X})$ that predicts the functional label $y\!\in\!\mathbf{Y}$ of an unseen data-specific graph.
  • Task-2: Targeted subgraph explanation. Develop an explainer $E(\cdot)$ that identifies the smallest subgraph $\hat{\mathcal{G}}\!\subseteq\!\mathcal{G}$ such that $F(\hat{\mathcal{G}},\mathbf{X})$ still outputs $y$.
Problem 2 (Pathway modeling). Many biological functions arise from long, multi-step reaction pathways that span several hops in $\mathcal{G}$. Capturing long-range dependencies is essential: (1) for functional prediction, as perturbation effects often propagate across distant nodes; and (2) for mechanistic insight into causal pathways beyond local interactions. Different experimental conditions on the same network, therefore yield distinct, data-specific subgraphs $\hat{\mathcal{G}}$, each revealing the targeted pathway most responsible for the given data. Hence,
  • Expectation: Pathway-level encoding and explaining.  The classifier $F(\cdot)$ and explainer $E(\cdot)$ leverage both graph topology and node features. $F(\cdot)$ aims to capture long-range dependencies, yielding high and class-balanced accuracy. Also, $E(\cdot)$ extracts subgraphs that retain biologically meaningful information of long pathways.
Proposed method
Framework
@@
ExPath comprises two components: graph-based classification and post-hoc subgraph explanation, as shown in Figure 2. To tackle Task-1, PathMamba, a classifier combining GNNs with state-space sequence modeling, is to capture both local node-pair interactions and global pathway-level dependencies. To address Task-2, PathExplainer, a graph explainer with pathway-wise masking, aims to identify the most influential subgraphs. We explicitly integrate pathway information into both models to meet Expectation.
Notably, our ExPath is compatible with large biological foundation models for encoding experimental data. In this work, we leverage large protein language model encodings to investigate the mapping from amino acid (AA) sequences to corresponding pathway bio-networks. Learning from AA sequence data is challenging due to its inherent complexity. Even slight variations can lead to significant structural changes, potentially disrupting protein functionality within pathways. Several studies focus on feature extraction in AA sequences, like AlphaFold (Jumper et al. 2021).
Feedforward Process. We first encode experimental data into node attributes using ESM-2 (Lin et al. 2023). It is pre-trained on over 60 million AA sequences with parameter scaling up to 15 billion. We then train PathMamba to learn pathway-level information and perform bio-network classification. Finally, we apply PathExplainer to selectively highlight the minimal subgraphs that drive the final prediction, offering interpretable insights into key pathways.
PathMamba: Pathway Representation Learning
@@
PathMamba integrates the Graph Isomorphism Network (GIN) with a novel pathway-wise Mamba model. It leverages the strengths of both global selective modeling mechanisms and message-passing GNNs. Specifically, inspired by GPS (Rampášek et al. 2022), our model avoids early-stage information loss that could arise from using GNNs in the initial layers. We employ novel pathway-wise global aggregation in efficient combination with random pathway sampling and sequential Mamba (Gu and Dao 2024) modeling.
Positional Encoding.
To address a fundamental limitation of GNNs (Xu et al. 2019) or hybrid models (Rampášek et al. 2022) to distinguish nodes with identical local structures, The node embedding $\mathbf{h}_{i}\in\mathbb{R}^{d}$ and the positional encoding $\mathbf{p}_{i}\in\mathbb{R}^{K}$ are concatenated and passed through a linear layer to obtain the final representation: $\mathbf{x}_{i}=\text{Linear}([\mathbf{h}_{i}\|\mathbf{p}_{i}]),$ where $\mathbf{h}_{i}\|\mathbf{p}_{i}]\in\mathbb{R}^{d+K}$ denotes the concatenation.
At each layer, node and edge features are updated by aggregating the outputs of a pathway-wise Mamba as:
$X^{l+1},$ $=$ $\texttt{PathMamba}^{l}\left(X^{l},A\right),$ (1)
$\textrm{computed as}\ \ \ X^{l+1}_{L},$ $=$ $\texttt{LocalGIN}^{l}\left(X^{l},A\right),$ (2)
$X^{l+1}_{G},$ $=$ $\texttt{GlobalMamba}^{l}\left(X^{l},A\right),$ (3)
$X^{l+1},$ $=$ $\texttt{MLP}^{l}\left(X^{l+1}_{L}+X^{l+1}_{G}\right),$ (4)
@@
Node-wise local aggregation.
The GINs update Node features by aggregating information from their local neighbors. The GIN operation can be expressed as:
$X^{l+1}_{L}=\text{ReLU}\left(W^{l}\cdot\big{(}(1+\epsilon)X^{l}+\sum_{j\in\mathcal{N}(i)}X^{l}_{j}\big{)}\right),$ (5)
where $\mathcal{N}(i)$ represents the set of neighbors of node $i$, $W^{l}$ is the learnable weight matrix at layer $l$, and $\epsilon$ is a trainable parameter controlling the importance of self-loops.
Pathway-wise global aggregation.
To capture long-term dependencies, we propose random pathway sampling and sequential pathway modeling in PathMamba.